# A Survey on Cross-Domain Sequential Recommendation

### 10 Jan 2024 v1

@@ Abstract
Cross-domain sequential recommendation (CDSR) shifts the modeling of user preferences from flat to stereoscopic by integrating and learning interaction information from multiple domains at different granularities (ranging from inter-sequence to intra-sequence and from single-domain to cross-domain). In this survey, we initially define the CDSR problem using a four-dimensional tensor and then analyze its multi-type input representations under multidirectional dimensionality reductions. Following that, we provide a systematic overview from both macro and micro views. From a macro view, we abstract the multi-level fusion structures of various models across domains and discuss their bridges for fusion. From a micro view, focusing on the existing models, we specifically discuss the basic technologies and then explain the auxiliary learning technologies. Finally, we exhibit the available public datasets and the representative experimental results as well as provide some insights into future directions for research in CDSR.
@@ 1 Introduction
In real life, users constantly leave interaction traces in multiple scenarios (e.g., shopping scenarios, reading scenarios, etc.), multiple platforms (e.g., Amazon, Taobao, etc.), and even multiple boards (e.g., books and movies in Douban, etc.). Considering all these as different domains, combining information across domains and shifting the modeling of user preferences from flat to stereoscopic is a new and important direction for the development of recommender systems. As shown in Table 1, starting from traditional one-class collaborative filtering (OCCF), further progressing to sequential one-class collaborative filtering (SOCCF) and to cross-domain one-class collaborative filtering (CD-OCCF), finally to cross-domain sequential one-class collaborative filtering (CD-SOCCF, also called cross-domain sequential recommendation, CDSR), recommender systems focus on information that is becoming deeper and more comprehensive.
@@
In the case of OCCF Rendle et al. (2012), the available information is limited to whether the user interacted with the item. As illustrated in Figure 1, we only know the items that both user 1 and user 2 have interacted with in a single domain, without considering the sequential pattern among successive items. SOCCF Lin et al. (2020), building upon OCCF, places greater emphasis on the order relationships between items within a sequence, i.e., the information of intra-sequence in a single domain, and thus the model can capture the user’s long-term interests and short-term preferences. At a large granularity, expanding from a single domain to multiple domains offers a fresh perspective to rich information for recommendation. As shown in Figure 1, users have interactions not only in the book domain but also in the movie domain and they are closely related. It constitutes CD-OCCF Hu et al. (2018) which integrates the interactions of both two domains. In this cross-domain scenario, designing models to transfer or aggregate interaction records from multiple domains, as well as aligning representations of users or items across different domains, are critical challenges.
Indeed, CDSR combines all of the directions aforementioned and further incorporates sequential information based on CD-OCCF and cross-domain information based on SOCCF. As indicated by the purple arrows in Figure 1, CDSR especially necessitates the precise capturing of sequential relationships between items across different domains. For example, user 1, after reading two books in the $Harry\ Potter$ series, chooses to watch the corresponding $Harry\ Potter$ movie. On the other hand, user 2 who enjoys watching romantic movies also reads the original novels after watching the $Pride\ and\ Prejudice$ movie. In addition to capturing the cross-domain sequential dependencies described above as much as possible, the challenges of CDSR include how to fuse information from different domains and how to distinguish between users’ specific preferences within a single domain and global preferences shared across multiple domains.
In this survey, we initially provide an overarching problem definition and modeling of CDSR, considering various dimensionality reductions and input representations. Then we adopt both a macro and micro view to summarize the existing works in CDSR. From a macro view, we present an overview of multi-level fusion structures, discussing how information is fused across different domains and exploring bridges for cross-domain fusion. From a micro view, we conduct a detailed analysis of various technologies employed by existing models that are categorized into basic technologies and auxiliary learning. Based on this analysis, we synthesize a comprehensive framework that provides an overview of the key technologies used in CDSR models. Furthermore, we also list the datasets commonly used in CDSR and the representative experimental results as well as provide some insights into potential future directions.
@@ 2 Formulation
In this section, we begin by presenting the problem definition that unifies CDSR using a four-dimensional tensor. Following that, we delve into the various directions of dimensionality reduction employed in practical modeling, as illustrated in Figure 2, along with the discussion of multi-type input representations.
@@ 2.1 Problem Definition
In a typical cross-domain sequential recommendation scenario, as temporal information and cross-domain information are incorporated, the interaction data between users and items undergoes a gradual expansion into a four-dimensional data tensor, consisting of dimensions pertaining to users, items, time, and domains. We denote this four-dimensional data tensor with $\Gamma\in\mathbb{R}^{n\times m\times s\times k}$, where $n$ is the number of users, $m$ is the number of items, $s$ is the discrete time intervals, and $k$ is the number of domains. Each element $\gamma_{(u,i,t,d)}$ in it signifies whether there is an interaction between user $u$ and item $i$ at time $t$ in domain $d$. Then the goal of CDSR is to estimate the probability for all the candidate items in each domain and to recommend the most probable next item to each user. The estimated probability can be formalized as follows,
@@
where $\hat{i}$ is the candidate item in each domain, $\Gamma$ is the four-dimensional tensor containing all interaction information, and $f(\ \Gamma\ )$ is the learned function to estimate $P(\ \hat{i}\ |\ \Gamma\ )$.
In fact, most existing works in CDSR are primarily established on two domains, i.e., $k=2$. Therefore, for the subsequent discussion, we take two domains as an example, representing them as domain A and domain B, respectively. Some researchers Bi et al. (2020); Xu et al. (2023c) also refer to the two domains as the source domain and the target domain. Moreover, some works Ma et al. (2019); Sun et al. (2023); Guo et al. (2021, 2023c) introduce an additional assumption with a shared account. They assume that within a real account, there are $q$ virtual users ($u_{j},j\leq q$) simultaneously active. Regarding the shared account issue, it is applicable not only in cross-domain scenarios but also in single-domain scenarios.
@@ 2.2 Multidirectional Dimensionality Reduction
For modeling and computational convenience, the four-dimensional tensor is often reduced dimensionally into corresponding three-dimensional tensors from various directions, as shown in Figure 2. In the following, we offer explanations for the dimensionality reduction from three different directions.
  • W/o User Dimension. In the modeling of CDSR, focusing on a user can give rise to a cross-domain interaction sequence, which can then serve as the basis for constructing a local graph in subsequent work.
  • W/o Time Dimension. Some involve mapping the four-dimensional tensor along the temporal dimension and aggregating all the interactions within a time, laying the groundwork for building a global graph in subsequent steps.
  • W/o Domain Dimension. Furthermore, we blend two or more domains into a hybrid domain, encompassing users and items from all domains, along with their interactions recorded in chronological order. Many existing studies treat the hybrid domain as an independent domain, designing model fusion structures that incorporate all three domains in parallel, which is shown in Section 3.1.
Indeed, these dimensionality reduction directions are often combined and used in conjunction rather than being treated as independent. Next, we visualize those dimensionality reduction processes on different types of inputs.
@@ 2.3 Multi-Type Input Representations
Researchers’ considerations on dimensionality reduction directions in CDSR are first reflected in the input representations. Here, we categorize conventional input representations into pure sequential representation and graph-encoded sequential representation. Subsequently, we analyze unconventional inputs, including side information and features obtained through pre-training.
@@ 2.3.1 Pure Sequential Representation
Arranging the interacted items of a user in chronological order yields a sequence. For each user $u$, we define $S^{A}_{u}=\{i^{A}_{1},i^{A}_{2},\dots\}$ and $S^{B}_{u}=\{i^{B}_{1},i^{B}_{2},\dots\}$ as his/her interaction sequences in domain A and domain B, respectively. If $u$ is an overlapping user who has interactions both in domain A and domain B, we can mix $S^{A}_{u}$ and $S^{B}_{u}$ to a hybrid sequence in chronological order (e.g., $S^{hybrid}_{u}=\{i^{A}_{1},i^{B}_{1},i^{A}_{2},i^{A}_{3},i^{B}_{2},\dots\}$), and view it as the hybrid domain. From a pure sequential perspective, we can substitute the four-dimensional tensor in Eq. (1) with the set of sequences in each domain, as follows,
@@ 2.3.3 Side Information
More and more researchers consider incorporating side information to enrich the semantic understanding of users’ historical behaviors. Here, we divide the most common side information into three categories: time, text (i.e., user/item context and review), and knowledge graph.
Time. Previously the timestamps are only used to order the items, but some researchers Xiao et al. (2023); Wang et al. (2022); Guo et al. (2022) define $t_{ij}=|t_{i}-t_{j}|$ to model the time interval and apply it to the subsequent learning of model, where $t_{i}$ and $t_{j}$ are the timestamps that the item $i$ and the item $j$ are interacted with, respectively. In addition to the time intervals, we can mine more information about time, such as periodicity and the duration of the interaction, etc.
User/Item Context and Review. Except for the user ID, triples composed of multiple contextual information (e.g., (``$UserID",``City",``Age",\cdots)$) are utilized as a basis for finding user-user relationships Ouyang et al. (2020). Similarly, the context information that consists of categories, tags, keywords, etc, is used as a supplement to item ID Xiao et al. (2023); Ouyang et al. (2020); Zhuang et al. (2020). And review is also a type of available text message that can associate users and items. A common way to encode these texts is through some pre-trained models.
Knowledge Graph. Knowledge graphs Bi et al. (2020); Ma et al. (2022) deal with information from both entity and relationship views. It not only contains structural information between nodes but also implies some relationship description. A knowledge graph defines an entity set $E^{KG}$ and a relation set $R^{KG}$, which consists of multiple entity-relation-entity triples $<e_{i},r,e_{j}>$ (e.g., $<e_{a_{1}},Is\_the\_same\_category,e_{b_{1}}>$ meaning that the entity $e_{a_{1}}$ from domain A has the same category as entity $e_{b_{1}}$ from domain B).
@@ 2.3.4 Pre-trained Features
Considering the privacy issues in two or more domains, some models Ding et al. (2023); Zhang et al. (2023a); Lei et al. (2021) use model-trained features as the input of the source domain, instead of raw data. Ding et al. (2023) directly combines the pre-trained user features from one domain with another domain. Zhang et al. (2023a) utilizes federated learning to fetch the global representation from the server, and then injects it into the local model.
@@ 3 Macro-View: What Structure Is Used to Fuse the CDSR Information?
In CDSR, the fusion structure serves as the primary framework of the model, providing pathways for the separation and aggregation of these features. In this section, we first propose the multi-level fusion structures and then elaborate on the bridge for inter-domain fusion.
@@ 3.1.1 One-Level Fusion Structures
In earlier works, to combine the cross-domain information, there are a lot of methods that first learn user preferences from domain A and domain B separately, and then fuse the representations of the two domains using various operations (shown in Figure 3(a)). These operations are not limited to concatenation Lei et al. (2021); Guo et al. (2023c), summation Alharbi and Caragea (2021); Ding et al. (2023), multi-layer perceptron (MLP) Bi et al. (2020), or some attention mechanisms Ouyang et al. (2020); Li et al. (2021).
Additionally, some researchers employ transfer learning to transfer knowledge from a source domain to a target domain Chen et al. (2021); Liu and Zhu (2021), or train a discriminator to bridge the representation of two domains with the idea of adversarial learning Li et al. (2022), shown in Figure 3(b).
In contrast to the above-juxtaposed structures, there are also works utilizing a tandem structure to fuse information Alharbi and Caragea (2022) (i.e., Figure 3(c)), or tackling the problem from the perspective of a hybrid-domain view (i.e., Figure 3(d)). Some researchers Guo et al. (2021, 2022) construct a global graph for the hybrid domain to combine the cross-domain knowledge. And others like Ma et al. (2019); Sun et al. (2023) learn the item transition patterns in the hybrid sequences $S^{hybrid}$.
@@ 3.1.2 Two/Three-Level Fusion Structures
Considering the potential for further exploration and advancement in the field of fusion structure, researchers combine the single domain and the hybrid domain with a multi-level fusion structure. As shown in Figure 3(e), in order to predict the next item in a separate domain, the hybrid domain is utilized as the main sharer Cao et al. (2022); Wang et al. (2022) or the bridge Zheng et al. (2022) to transfer knowledge from another domain. Some researchers Ye et al. (2023) choose to combine the domain A and domain B first and then fuse the hybrid information (i.e., Figure 3(f)).
@@ 3.1.3 Discussion
Although multi-level fusion structures have been proposed, it does not mean that a more complex structure will result in a better model. The simple fusion structures, i.e., one-level fusion structures, are easy to implement but may not be conducive to comprehensive modeling of domain-specific and domain-generic features due to the limitation of the degree of fusion. While improving effectiveness, multi-level fusion structures bring increased complexity and reduced interpretability.
@@ 3.2 Bridges for Cross-Domain Fusion
When considering the overall cross-domain fusion structure, it is important to note the bridges through which information from different domains can be shared. We categorize those bridges of the existing works into three categories, i.e., the same users, contexts, and clusters/groups, as shown in Figure 4. The first one primarily exists in scenarios with overlapping users, while the latter two can be employed in situations without any overlapping users.
  • The Same Users. Building a bridge with overlapping users is the most common practice. Most works confine recommendation scenarios to fully overlapping users, without considering any non-overlapping users. However, some works Li et al. (2022) consider both overlapping users and non-overlapping users.
  • Contexts. In the absence of overlapping users, leveraging semantic similarity in natural language can also serve as a bridge to connect items between different domains. For example, using item contexts instead of traditional item ID can more easily identify similarities between items Liu et al. (2023).
  • Clusters/Groups. If there are neither overlapping users nor side information available, the approach of using clusters/groups as a bridge can be applicable Lin et al. (2023a), due to the similarity in preferences within a certain group of people.
Different bridges exhibit distinct characteristics, each suitable for different scenarios. Methods relying on overlapping users often yield better performance but may face limitations due to sparse data in real-world scenarios. On the other hand, methods built on non-overlapping users can be applicable to a wider range of scenarios but the performance may be limited.

### 19 Jan 2024 v2

@@ Abstract
Cross-domain sequential recommendation (CDSR) shifts the modeling of user preferences from flat to stereoscopic by integrating and learning interaction information from multiple domains at different granularities (ranging from inter-sequence to intra-sequence and from single-domain to cross-domain). In this survey, we first define the CDSR problem using a four-dimensional tensor and then analyze its multi-type input representations under multidirectional dimensionality reductions. Following that, we provide a systematic overview from both macro and micro views. From a macro view, we abstract the multi-level fusion structures of various models across domains and discuss their bridges for fusion. From a micro view, focusing on the existing models, we specifically discuss the basic technologies and then explain the auxiliary learning technologies. Finally, we exhibit the available public datasets and the representative experimental results as well as provide some insights into future directions for research in CDSR.
@@ 1 Introduction
In real life, people leave interaction traces in multiple scenarios (e.g., shopping scenarios, reading scenarios, etc.), multiple platforms (e.g., Amazon, Taobao, etc.), and even multiple boards (e.g., books and movies in Douban, etc.) all the time. Considering all these as different domains, it is a critical step for the development of recommender systems to combine information across domains so that the data of user preferences is no longer fragmented, and shifting the modeling from flat to stereoscopic. As shown in Table 1, recommender systems focus on information that is becoming deeper and more comprehensive, starting from traditional one-class collaborative filtering (OCCF), further progressing to sequential one-class collaborative filtering (SOCCF) and to cross-domain one-class collaborative filtering (CD-OCCF), finally to cross-domain sequential one-class collaborative filtering (CD-SOCCF, also called cross-domain sequential recommendation, CDSR).
@@
In the case of OCCF Rendle et al. (2012), the available information is limited to whether the user interacted with the item. As illustrated in Figure 1, we only know all the items that user 1 and user 2 have interacted with in a single domain, without considering the sequential pattern among successive items. SOCCF Lin et al. (2020), building upon OCCF, places greater emphasis on the order relationships between items within a sequence, i.e., the information of intra-sequence in a single domain, and thus the model can capture the user’s long-term and short-term preferences. At a large granularity, expanding from a single domain to multiple domains offers a fresh perspective to address sparse data issues in recommender systems, thus constituting CD-OCCF Hu et al. (2018). For instance, as shown in Figure 1, users have interactions not only in the book domain but also in the closely related movie domain. In the cross-domain scenario, we must meet the challenges that transferring or aggregating interaction records from multiple domains as well as aligning representations of users or items across different domains.
Indeed, CDSR combines all of the aforementioned directions and further incorporates sequential information based on CD-OCCF and cross-domain information based on SOCCF. As indicated by the purple arrows in Figure 1, CDSR necessitates the precise capturing of sequential relationships between items across different domains. For example, user 1, after reading two books in the $Harry\ Potter$ series, chooses to watch the corresponding $Harry\ Potter$ movie. Similarly, user 2 who enjoys watching romantic movies also reads the original novels after watching the $Pride\ and\ Prejudice$ movie. Besides, it is also worth considering in CDSR how to fuse sequential information from different domains as well as how to distinguish between users’ specific preferences within a single domain and global preferences shared across multiple domains.
In this survey, we first formulate CDSR’s problem definition and modeling, considering various dimensionality reductions and different input representations. Then we adopt both a macro and micro view to summarize the existing works in CDSR. From a macro view, we present an overview of multi-level fusion structures, discussing how to fuse information across different domains and exploring bridges for cross-domain fusion. From a micro view, we conduct a detailed analysis of various technologies employed by existing works that are categorized into basic technologies and auxiliary learning. Based on this analysis, we synthesize a comprehensive classification table (as shown in Table 2) and framework figure of the key technologies (as shown in Figure 5). Furthermore, we also list the datasets commonly used in CDSR and the representative experimental results as well as provide some insights into potential future directions.
@@ 2 Formulation
In this section, we first present the problem definition that unifies CDSR using a four-dimensional tensor. Then, referring to Figure 2, we delve into the various directions of dimensionality reduction that are often applied in practical modeling and the corresponding multi-type input representations.
@@ 2.1 Problem Definition
As temporal information and cross-domain information are incorporated, the interaction data between users and items in CDSR undergoes a gradual expansion into a four-dimensional data tensor, consisting of dimensions about users, items, time, and domains. We denote this four-dimensional tensor with $\Gamma\in\mathbb{R}^{n\times m\times s\times k}$, where $n$ is the number of users, $m$ is the number of items, $s$ is the discrete time intervals, and $k$ is the number of domains. Each element $\gamma_{(u,i,t,d)}$ in it signifies whether there is an interaction between user $u$ and item $i$ at time $t$ in domain $d$. The goal of CDSR is to estimate the probability for all candidate items in each domain, and then to recommend the most probable next item to each user. The estimated probability can be formalized as follows,
@@
where $\hat{i}$ denotes the candidate item in each domain, $\Gamma$ is the four-dimensional tensor containing all interaction information, and $f(\ \Gamma\ )$ indicates the learned function to estimate $P(\ \hat{i}\ |\ \Gamma\ )$.
Notice that most existing works in CDSR are primarily established on two domains, i.e., $k=2$. Therefore, for the subsequent discussion, we take two domains as an example and represent them as domain A and domain B, respectively. In fact, some researchers Bi et al. (2020); Xu et al. (2023c) also refer to the two domains as the source domain and the target domain. Moreover, some works Ma et al. (2019); Sun et al. (2023); Guo et al. (2021, 2023c) introduce an additional assumption with a shared account. They assume that within a real account, there are $q$ virtual users ($u_{j},j\leq q$) simultaneously active. Regarding the shared account issue, it is applicable not only in cross-domain scenarios but also in single-domain scenarios.
@@ 2.2 Multidirectional Dimensionality Reduction
For computational and modeling convenience, the four-dimensional tensor is often reduced to three-dimensional tensors from various directions (as shown in Figure 2). Below, we explain dimensionality reduction from three directions:
  • W/o User Dimension. For each user, their interactions in chronological order can form a sequence in each domain. These sequences reflect each user’s personalization as well as long and short-term preferences in different domains, and can further serve as the basis for constructing a local graph in subsequent work.
  • W/o Time Dimension. Mapping the tensor along the temporal dimension and aggregating interactions within a time lays the groundwork for a global graph. It allows models to learn user-to-user and user-to-item connections, which are more skewed toward users’ global preferences.
  • W/o Domain Dimension. Blending two or more domains into a hybrid domain, so that the models can clarify the passing relationships between items originally in different domains. Many existing studies treat the hybrid domain as an independent domain, designing model fusion structures that incorporate all three domains in parallel, which is shown in Section 3.1.
Indeed, these dimensionality reduction directions are often combined and used in conjunction rather than being treated as independent. Next, we visualize these reduction processes on different input representations.
@@ 2.3 Multi-Type Input Representations
Researchers’ considerations on dimensionality reduction directions in CDSR are first reflected in the input representations. Here, we categorize conventional input representations into pure sequential representation and graph-encoded sequential representation. Subsequently, we analyze unconventional inputs, including side information and pre-trained features.
@@ 2.3.1 Pure Sequential Representation
Arranging the interacted items of a user in chronological order yields a sequence. For each user $u$, we define $S^{A}_{u}=\{i^{A}_{1},i^{A}_{2},\dots\}$ and $S^{B}_{u}=\{i^{B}_{1},i^{B}_{2},\dots\}$ as his/her interaction sequences in domain A and domain B, respectively. If $u$ is an overlapping user who has interactions both in domain A and domain B, we can mix $S^{A}_{u}$ and $S^{B}_{u}$ to a hybrid sequence in chronological order (e.g., $S^{hybrid}_{u}=\{i^{A}_{1},i^{B}_{1},i^{A}_{2},i^{A}_{3},i^{B}_{2},\dots\}$), which is viewed as a sequence of the hybrid domain. From a pure sequential perspective, we can substitute the four-dimensional tensor in Eq. (1) with the set of sequences in each domain, as follows,
@@ 2.3.3 Side Information
More and more researchers consider incorporating side information to enrich the semantic representation of users’ historical behaviors. Here, we divide the most common side information into three categories: time, text (i.e., user/item context and review), and knowledge graph.
Time. Previously the timestamps are only used to order the items, but some researchers Xiao et al. (2023); Wang et al. (2022); Guo et al. (2022) define $t_{ij}=|t_{i}-t_{j}|$ to model the time interval and apply it to the subsequent learning of model, where $t_{i}$ and $t_{j}$ are the timestamps that the item $i$ and the item $j$ are interacted with, respectively. Besides, we can also mine more information about time, such as periodicity and the duration of the interaction, etc.
User/Item Context and Review. Except for the user ID, triples composed of multiple contextual information (e.g., (``$UserID",``City",``Age",\cdots)$) are utilized as a basis for finding user-user relationships Ouyang et al. (2020). Similarly, the context information that consists of categories, tags, keywords, etc, is used as a supplement to item ID Xiao et al. (2023); Ouyang et al. (2020); Zhuang et al. (2020). Review as a type of available text message can associate users and items well. A common way to encode these texts is through some pre-trained models.
Knowledge Graph. Knowledge graphs Bi et al. (2020); Ma et al. (2022) deal with information from both entity and relationship views. It not only contains structural information between nodes but also implies some relationship description. A knowledge graph defines an entity set $E^{KG}$ and a relation set $R^{KG}$, which consists of multiple entity-relation-entity triples $<e_{i},r,e_{j}>$ (e.g., $<e_{i_{1}^{A}},Is\_the\_same\_category,e_{i_{1}^{B}}>$ meaning that the entity $e_{i_{1}^{A}}$ from domain A has the same category as entity $e_{i_{1}^{B}}$ from domain B).
@@ 2.3.4 Pre-trained Features
Considering the privacy issues in two or more domains, some works Ding et al. (2023); Zhang et al. (2023a); Lei et al. (2021) use model-trained features as the input of the source domain, instead of raw data. Ding et al. (2023) directly combines the pre-trained user features from one domain with another domain. Zhang et al. (2023a) utilizes federated learning to fetch the global representation from the server, and then injects it into the local model.
@@ 3 Macro-View: What Structure Is Used to Fuse the CDSR Information?
In CDSR, the fusion structure serves as the primary skeleton of the model, providing pathways for the separation and aggregation of these features. In this section, we first propose the multi-level fusion structures and then elaborate on the bridge for inter-domain fusion.
@@ 3.1.1 One-Level Fusion Structures
In earlier works, to combine the cross-domain information, there are a lot of works that first learn user preferences from domain A and domain B separately, and then fuse the representations of the two domains using various operations (shown in Figure 3(a)). These operations are not limited to concatenation Lei et al. (2021); Guo et al. (2023c), summation Alharbi and Caragea (2021); Ding et al. (2023), multi-layer perceptron (MLP) Bi et al. (2020), or some attention mechanisms Ouyang et al. (2020); Li et al. (2021).
Additionally, some works employ transfer learning to transfer knowledge from a source domain to a target domain Chen et al. (2021); Liu and Zhu (2021), or train a discriminator to bridge the representation of two domains with the idea of adversarial learning Li et al. (2022), shown in Figure 3(b).
In contrast to the above-juxtaposed structures, there are also works utilizing a tandem structure to fuse information Alharbi and Caragea (2022) (i.e., Figure 3(c)), or directly tackling the problem from the perspective of a hybrid-domain view (i.e., Figure 3(d)). For instance, some works Guo et al. (2021, 2022) construct a global graph for the hybrid domain to combine the cross-domain knowledge. And others like Ma et al. (2019); Sun et al. (2023) choose to learn the item transition patterns in the hybrid sequences $S^{hybrid}$.
@@ 3.1.2 Two/Three-Level Fusion Structures
Considering the potential for further exploration and advancement in the field of fusion structure, some works combine the single domain and the hybrid domain with a multi-level fusion structure. As shown in Figure 3(e), in order to predict the next item in a separate domain, the hybrid domain is utilized as the main sharer Cao et al. (2022); Wang et al. (2022) or the bridge Zheng et al. (2022) to transfer knowledge from another domain. There are also some works Ye et al. (2023) that choose to combine the domain A and domain B first and then fuse the hybrid information (i.e., Figure 3(f)).
@@ 3.1.3 Discussion
Although we propose multi-level fusion structures, it does not mean that a more complex structure will result in a better performance. The simple fusion structures, i.e., one-level fusion structures, are easy to implement but may not be conducive to comprehensive modeling of domain-specific and domain-generic features due to the limitation of the degree of fusion. While improving effectiveness, multi-level fusion structures bring increased complexity and reduced interpretability.
@@ 3.2 Bridges for Cross-Domain Fusion
When comes to cross-domain fusion, it’s vital to clarify the bridges for information sharing between domains. We divide those bridges into three categories: the same users, contexts, and clusters/groups, as shown in Figure 4. The first is prevalent in scenarios with overlapping users, while the latter two apply in scenarios without overlapping users.
  • The Same Users. If the same users exist, then they are the first choice as the bridge. Some works only focus on fully overlapping users, neglecting non-overlapping users, while others Li et al. (2022) consider both cases.
  • Contexts. If the same users don’t exist, it is feasible to leverage the semantic similarity of natural language. For instance, we can replace traditional item IDs with item contexts to explore similarities between items of different domains Liu et al. (2023).
  • Clusters/Groups. Further restricting the condition to no overlapping users and no side information, we can work from the clusters/groups, since a specific group of users could have similar preferences Lin et al. (2023a).
Different bridges exhibit distinct characteristics, each suitable for different scenarios. Works relying on overlapping users often perform well but their application scenarios may face limitations due to sparse real-world data. Conversely, works built on non-overlapping users can be applied to a broader range of scenarios but may have limited performance.